# Multiple Queries with Multiple Keys: A Precise Prompt Matching Paradigm for Prompt-based Continual Learning

### v2

@@ 5.4 Setting Fine-grained Keys
Figure 6: t-SNE (Van der Maaten & Hinton, 2008) visualization of the Queries. The samples come from three classes of three tasks in the 5-task Split Imagenet-R. The shapes represent the sample tasks, and the colors represent the sample categories.
5.5 Going Deeper and Longer
As shown in Figure 5, the performance of the MQMK method improves as the e-prompt depth and length increase, reaching its optimal performance of 78.82% when the e-prompt depth is 10 and length is 80. In contrast, the performance of the SQSK method decreases as the prompts go deeper and longer. A possible reason is that increasing the depth and length of the prompt can improve its quality. The higher matching rate of MQMK allows it to effectively utilize the higher-quality prompt, thereby enhancing performance. In contrast, SQSK does not benefit in the same way. This suggests that an accurate query paradigm can allow the prompt to go deeper and longer, thereby improving model performance.
5.6 Visualization of Queries
We present visualizations of the queries selected by MQ, as well as queries in SQ, as shown in Figure 6. It can be observed that the queries selected by MQ exhibit stronger intra-class cohesion and inter-class separation, indicating that MQ queries are precise and the introduction of prompts is necessary. On the contrary, due to SQ’s reliance solely on ViT’s generalization ability and the lack of task-specific knowledge, the queries of SQ are highly dispersed, resulting in almost failed clustering. Additionally, although the queries show a tendency to cluster by task, the smallest granularity of clustering is at the class level, which indicates the necessity of setting class-level keys.
@@ 5.6 Visualization of Queries
Figure 7: The 3D visualization of specific prompt selection process of MQMK and SQSK on 10-task Split Imagenet-R.
Dataset SQSK MQMK
ImageNet-R 45.15 77.97
CIFAR-100 49.14 82.28
DomainNet 70.55 82.24
Table 5: Matching rate (%) on three benchmarks.
5.7 Prompt Matching Comparison
@@
As shown in Table 5, MQMK improves the matching rate by over 30% compared to SQSK on the challenging CIFAR100 and ImageNet-R. Due to the significant differences in data styles across different domains, SQSK can achieve high performance on DomainNet. To gain a deeper understanding of the prompt selection mechanism, we visualize the specific prompt selection process of MQMK and SQSK, generating two 3D bar charts, as shown in Figure 7. The results indicate that MQMK exhibits a significantly higher probability of selecting the true prompt compared to SQSK.
6 Conclusion
@@
To address the issue of low prompt selection accuracy in prompt-based continual learning, we propose the Multiple Queries with Multiple Keys prompt matching paradigm. Multiple Queries achieve precise feature extraction by incorporating task-specific knowledge. Multiple Keys, through fine-grained learnable keys, better represent the feature distribution of the training samples. They complement each other, significantly improving the matching rate and performance. However, Multiple Queries introduce a higher inference burden, and exploring ways to accelerate the inference phase with Multiple Queries is a promising direction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Key Matching Classifier
@@
Table 6: $A_{T}$ (%) of MQMK using 3 different classifiers on 10-task Split Imagenet-R.
According to Equations 2 and 3, when $K=1$, our method can be interpreted as selecting the key with the highest cosine similarity, and choosing the corresponding prompt. Since MK extends keys to the class level, establishing an identity mapping between keys and classes and directly deriving the classification result upon key selection is more straightforward and worth exploring. We refer to this classifier, which directly classifies based on the selected class-level keys, as the Key Matching (KM) classifier.
We compare the performance of the KM classifier with the Fully Connected (FC) classifier used throughout the experiment and an additional classic prototype-based Nearest Class Mean (NCM) classifier, as shown in Table 6. The performance of the KM classifier was not satisfactory, failing to surpass both the FC and NCM classifiers. Notably, on the 10-task Split ImageNet-R, the matching rate reached 77.97%, as depicted in Table 5. Since the classification accuracy of the KM classifier essentially reflects the accuracy of key matching, while the matching rate represents the accuracy of prompt selection, the gap between them (5.70%) corresponds to the samples where key matching errors occur but the prompt is correctly selected. The matching rate can be considered a loose upper bound of the KM classifier’s accuracy. In this task, the matching rate is lower than the FC classifier’s accuracy, which explains why the KM classifier fails to outperform the FC classifier. However, if the matching rate were to improve further, it remains to be explored whether the KM classifier could achieve performance comparable to or even surpassing that of the FC classifier.
Appendix B Discussion on Computational Costs and Number of Parameters
Method Training (GFLOPs) Inference (GFLOPs) Learnable Parameters (M) All Parameters (M)
SQSK 46.92 33.22 5.63 87.88
MQMK 30.56 164.10 5.70 87.95
Table 7: Computational costs and number of parameters of SQSK and MQMK with $L_{e}=40,H_{e}=6,L_{g}=5,H_{g}=2$ on 10-task Split CIFAR-100 dataset.
In Section 4.4, we have discussed the theoretical inference cost of MQMK. In this section, we will discuss the practical computational costs, theoretical training computational costs, and the number of parameters.
@@ Appendix B Discussion on Computational Costs and Number of Parameters
Training Process. MQ does not need to be executed. Instead, queries are directly located based on the task $t$. Since the prediction and the query use the same features, only one forward and one backward pass through the ViT backbone is required. In contrast, SQSK needs to perform both a query and a prediction with the prompt, which requires two forward passes and one backward pass through the ViT backbone.
Practical Results. As shown in Table 7, during the training phase, the computational cost of MQMK is 0.65 times that of SQSK. This is because MQ performs one less forward pass of the ViT. In the inference phase, MQMK’s computational cost is approximately five times that of SQSK. This is due to the multiple forward passes of ViT caused by MQ. The parameter count of MQMK is nearly the same as that of SQSK. Since the number of parameters for the keys is relatively small compared to the prompt, the additional parameter burden introduced by MK is not significant. The practical results align with the theoretical expectations.
Appendix C Visualization of Features in Perfect Match
@@
Figure 8: t-SNE visualization of features in perfect match. The samples come from two classes across five tasks in the 5-task Split Imagenet-R. The shapes represent the sample tasks, and the colors represent the sample categories.
@@
In Section 5.6, we have discussed the query visualization of MQ and SQ. In this section, we provide feature visualizations under the condition where the model selects the true prompt for all samples (perfect match), which can be considered an ideal upper bound. As shown in Figure 8, when the true prompt is selected, the discriminability of the features is very strong. Therefore, improving prompt matching rate can significantly enhance performance. Additionally, there is no inherent relationship between classes within the same task. However, since the same prompt is selected, the features of samples in the same task become very similar. It suggests that different prompts generate different subspaces, and selecting the wrong prompt may cause subspace drift, leading to biased predictions. This highlights that the selection of the prompt is crucial.
Appendix D Performance of True and False Prompt in MQMK
@@
Figure 9: The average accuracy for four scenarios: when MQMK selects False Prompt and True Prompt, when samples initially with False Prompt are manually replaced with True Prompt, and when all samples use True Prompt.
@@
In Section 3.3, we have already discussed the average accuracy of the samples selected with True and False Prompts in SQSK. In this section, we will discuss the situation in MQMK and compare the two scenarios. Compared to SQSK, MQMK has a lower accuracy for the incorrectly selected False Prompt samples. Even when provided with the True Prompt, the performance remains only at 68.34%. This suggests that MQMK has already matched the more recognizable and straightforward samples, and the remaining samples are inherently difficult to recognize. Even with the True Prompt, the performance is limited. The samples selected with the True Prompt in MQMK are easier to match and recognize, achieving an accuracy of 89.01%.
Appendix E Effect of Top-$K$ in Category Aggregation
$K$ $A_{T}$ ($\uparrow$) Matching Rate (($\uparrow$( ↑)
@@ Appendix E Effect of Top-$K$ in Category Aggregation
Table 8: Average accuracy (%) and matching rate (%) under different $K$-category aggregation on 10-task Split Imagenet-R.
In previous experiments, we set $K$ to 1 at all times. Here, we will discuss the performance of aggregating multiple classes and then selecting the prompt. As shown in Table 8, when $K$ is 1, both performance and matching reach their optimal values. As $K$ increases, both performance and matching rate decline. This suggests that the classes within the same task are not similar in features, and aggregating the matching scores for a group of classes has a negative effect.
Appendix F Dataset Introduction
@@
  • CIFAR-100 is a popular dataset in machine learning, consisting of 100 different classes. It contains 60,000 32$\times$32 color images, divided into 50,000 training images and 10,000 test images. Each class contains 600 images, and the categories are diverse, such as animals, vehicles, and household items.

### v3

@@ 5.4. Setting Fine-grained Keys
Figure 6. t-SNE (Van der Maaten and Hinton, 2008) visualization of the Queries. The samples come from three classes of three tasks in the 5-task Split Imagenet-R. The shapes represent the sample tasks, and the colors represent the sample categories.
5.5. Going Deeper and Longer
As shown in fig. 5, the performance of the MQMK method improves as the e-prompt depth and length increase, reaching its optimal performance of 78.82% when the e-prompt depth is 10 and length is 80. In contrast, the performance of the SQSK method decreases as the prompts go deeper and longer. A possible reason is that increasing the depth and length of the prompt can improve its quality. The higher matching rate of MQMK allows it to effectively utilize the higher-quality prompt, thereby enhancing performance. In contrast, SQSK does not benefit in the same way. This suggests that an accurate query paradigm can allow the prompt to go deeper and longer, thereby improving model performance.
5.6. Visualization of Queries
We present visualizations of the queries selected by MQ, as well as queries in SQ, as shown in fig. 6. It can be observed that the queries selected by MQ exhibit stronger intra-class cohesion and inter-class separation, indicating that MQ queries are precise and the introduction of prompts is necessary. On the contrary, due to SQ’s reliance solely on ViT’s generalization ability and the lack of task-specific knowledge, the queries of SQ are highly dispersed, resulting in almost failed clustering. Additionally, although the queries show a tendency to cluster by task, the smallest granularity of clustering is at the class level, which indicates the necessity of setting class-level keys.
@@ 5.6. Visualization of Queries
Figure 7. The 3D visualization of specific prompt selection process of MQMK and SQSK on 10-task Split Imagenet-R.
6. Matching Rate Across Methods
@@
As shown in table 10, MQMK improves the matching rate by over 30% compared to SQSK on the challenging CIFAR-100 and ImageNet-R. Due to the significant differences in data styles across different domains, SQSK can achieve high performance on DomainNet. To gain a deeper understanding of the prompt selection mechanism, we visualize the specific prompt selection process of MQMK and SQSK, generating two 3D bar charts, as shown in fig. 7. The results indicate that MQMK exhibits a significantly higher probability of selecting the true prompt compared to SQSK. MQMK outperforms MQMK-EI on datasets with large style discrepancies, such as ImageNet-R and DomainNet, whereas on more consistent datasets like CIFAR-100, MQMK-EI, with its knowledge fusion enhanced queries, achieves higher matching rates. Both experiments demonstrate the effectiveness of our approach in improving the matching rate.
CODA-P and EvoPrompt use soft, weighted prompt selection, where correctness does not exist. The matching objective of ESN is task-specific classifiers, which bears similarity to prompt selection. According to ablation studies, MQMK is identical to other variants such as SQSK in all aspects except for the prompt strategy. In vertical comparisons with other methods, MQMK still demonstrates a clear advantage in terms of matching rate.
7. Conclusion
@@
To address the issue of low prompt selection accuracy in prompt-based continual learning, we propose the Multiple Queries with Multiple Keys prompt local matching paradigm. Multiple Queries achieve precise feature extraction by incorporating task-specific knowledge. Multiple Keys, through fine-grained learnable keys, better represent the feature distribution of the training samples. They complement each other, significantly improving the matching rate and performance through local matching mechanism. An accurate parameter selection strategy can bring significant performance improvements to existing continual learning methods based on parameter expansion and selection. Although precise querying incurs a certain query cost, MQMK-EI maintains high query accuracy while effectively reducing the query cost.
Acknowledgements.
This work was partially supported by the STI 2030-Major Projects of China (Grant No. 2021ZD0201300), the National Natural Science Foundation of China (Grant Nos. 62276127, 62495094), and the Fundamental Research Funds for the Central Universities (Grant No. 2024300394). The authors gratefully acknowledge these supports.
Supplementary Materials: Multiple Queries with Multiple Keys: A Precise Prompt Matching Paradigm for Prompt-based Continual Learning
1. Prompt Matching Rate
We begin by formally defining a True Prompt: for any test sample $(\mathbf{x}_{i},y_{i})$ drawn from task $t$, a prompt $p_{j}$ is considered true if and only if it was trained on task $t$, i.e., the sample’s data distribution matches the prompt’s training distribution ($I=t$).
Under the standard Single Query–Single Key (SQSK) paradigm, our experiments on Split ImageNet-R (10 tasks, 6,000 test samples) show that only 45.15% of samples are paired with their True Prompt.
To better analyze the impact of prompt selection errors, we divide the test set into two disjoint subsets based on whether SQSK correctly matches the True Prompt:
  • False Prompt: 1,232 samples for which SQSK selected an incorrect prompt.
  • True Prompt: 4,768 samples for which SQSK selected the correct prompt.
We then evaluate four experimental settings, shown in Figure 2 of the main text.
  • False Prompt (62.06%): Accuracy on the “False Prompt” subset when using the initially retrieved (incorrect) prompts.
  • True Prompt (79.33%): Accuracy on the “True Prompt” subset with correctly retrieved prompts.
  • False $\rightarrow$ True (88.68%): Accuracy on the same 1,232 “False Prompt” samples, after manually replacing the incorrect prompts with their corresponding True Prompts. Since the sample set remains unchanged, this directly measures the performance gain from correcting prompt selection.
  • Perfect Match (84.46%): Accuracy on the entire 6,000 test samples when each sample is matched with its True Prompt, representing an ideal upper bound.
Several important observations emerge from these results:
  1. (1)
    The comparison between ”False Prompt” and “False $\rightarrow$ True” is fair because they use exactly the same sample set; the large accuracy gain (+26.62%) demonstrates the significant negative impact of incorrect prompt selection.
  2. (2)
    Interestingly, “False $\rightarrow$ True” achieves even higher accuracy than the “True Prompt” subset. This may be because the samples initially mismatched by SQSK tend to belong to tasks that are harder to identify, but once the correct prompt is provided, these samples become relatively easier to classify within their task.
  3. (3)
    ”Perfect Match” represents the upper bound that could be achieved by an ideal prompt selection mechanism, confirming the importance of improving the prompt matching rate. It should be noted that perfect matching involves information leakage, so its accuracy cannot be used as a baseline for performance comparison, but only as a reference for the upper bound of performance.
In summary, these experiments clearly show that prompt selection is a critical factor affecting performance under SQSK. Motivated by this, we propose the Multiple Queries with Multiple Keys (MQMK) paradigm to improve the prompt matching rate and thus boost overall accuracy.
2. Key Matching Classifier
@@
Table 6. $A_{T}$ (%) of MQMK using 3 different classifiers on 10-task Split Imagenet-R.
When $K=1$, our method can be interpreted as selecting the key with the highest cosine similarity, and choosing the corresponding prompt. Since MK extends keys to the class level, establishing an identity mapping between keys and classes and directly deriving the classification result upon key selection is more straightforward and worth exploring. We refer to this classifier, which directly classifies based on the selected class-level keys, as the Key Matching (KM) classifier.
We compare the performance of the KM classifier with the Fully Connected (FC) classifier used throughout the experiment and an additional classic prototype-based Nearest Class Mean (NCM) classifier. As shown in table 6, the performance of the KM classifier was not satisfactory, failing to surpass both the FC and NCM classifiers. Notably, on the 10-task Split ImageNet-R, the matching rate reached 77.97%. Since the classification accuracy of the KM classifier essentially reflects the accuracy of key matching, while the matching rate represents the accuracy of prompt selection, the gap between them (5.70%) corresponds to the samples where key matching errors occur but the prompt is correctly selected. The matching rate can be considered a loose upper bound of the KM classifier’s accuracy. In this task, the matching rate is lower than the FC classifier’s accuracy, which explains why the KM classifier fails to outperform the FC classifier. However, if the matching rate were to improve further, it remains to be explored whether the KM classifier could achieve performance comparable to or even surpassing that of the FC classifier.
3. Discussion on Computational Costs and Number of Parameters
Method Training (h) Inference (ms) Learnable Parameters (M) All Parameters (M)
SQSK 4.28 14.65 5.63 87.88
MQMK 2.78 72.62 5.70 87.95
MQMK (Parallel) 2.78 7.48 5.70 87.95
MQMK-EI 2.78 14.86 5.70 87.95
Table 7. Computational costs and number of parameters for SQSK and MQMK with $L_{e}=40$, $H_{e}=7$, $L_{g}=5$, and $H_{g}=2$ on the 10-task split of CIFAR-100, evaluated on a single NVIDIA RTX 4090 GPU.
We have discussed the theoretical cost of MQMK. In this section, we will discuss the practical computational costs, theoretical training computational costs, and the number of parameters.
@@ 3. Discussion on Computational Costs and Number of Parameters
Training Process. Queries are directly located based on the task $t$. Since the prediction and the query use the same features, only one forward and one backward pass through the ViT backbone is required. In contrast, SQSK needs to perform both a query and a prediction with the prompt, which requires two forward passes and one backward pass through the ViT backbone.
Practical Results. Training Phase: The MQMK series is more efficient, saving approximately 1/3 of the training time. This is because SQ-based methods require running the ViT backbone twice—once for querying and once for inference using the selected prompt. In contrast, MQMK performs both querying and inference based on the same feature representation, avoiding redundant computation. Inference Phase: Both MQMK-EI and SQSK employ a two-stage sequential inference process, leading to comparable inference times. However, MQMK (Parallel) significantly improves inference efficiency by replicating each input $M$ times, combining them with different prompts, and forwarding them as a single batch of size $M$. This parallelization strategy results in nearly 2× speedup. While the sequential inference variant of MQMK incurs higher latency than other approaches, its speed is still acceptable in many real-world applications. The practical results align with the theoretical expectations.
4. Visualization of Features in Perfect Match
@@
Figure 8. t-SNE (Van der Maaten and Hinton, 2008) visualization of features in perfect match. The samples come from two classes across five tasks in the 5-task Split Imagenet-R. The shapes represent the sample tasks, and the colors represent the sample categories.
@@
We have discussed the query visualization of MQ and SQ. In this section, we provide feature visualizations under the condition where the model selects the true prompt for all samples (perfect match), which can be considered an ideal upper bound. As shown in fig. 8, when the true prompt is selected, the discriminability of the features is very strong. Therefore, improving prompt matching rate can significantly enhance performance. Additionally, there is no inherent relationship between classes within the same task. However, since the same prompt is selected, the features of samples in the same task become very similar. It suggests that different prompts generate different subspaces, and selecting the wrong prompt may cause subspace drift, leading to biased predictions. This highlights that the selection of the prompt is crucial.
Method 2 3 4 5 6 7 8 9 10
MQMK 86.53 86.27 86.13 85.73 85.20 84.53 84.40 84.40 83.73
MQMK-EI 84.93 84.80 84.93 84.33 84.66 83.73 83.73 82.67 83.33
Table 8. Accuracy of task 2 across time, on 10-task Split Imagenet-R
5. Visualization of Queries in MQMK-EI
(a) MQMK-EI.
(b) Selected Queries in MQ.
Figure 9. t-SNE visualization of the Queries. The samples come from three classes of three tasks in the 5-task Split Imagenet-R. The shapes represent the sample tasks, and the colors represent the sample categories.
As illustrated in fig. 9, the enhanced queries in MQMK-EI and the standard queries in MQMK both demonstrate improved separability compared to SQ. Notably, MQMK shows a slightly higher degree of query separability than MQMK-EI. These results indicate that MQMK-EI can be regarded as an efficient approximation to MQMK.
6. Performance of True and False Prompt in MQMK
@@
Figure 10. The average accuracy for four scenarios: when MQMK selects False Prompt and True Prompt, when samples initially with False Prompt are manually replaced with True Prompt, and when all samples use True Prompt.
@@
We have already discussed the average accuracy of the samples selected with True and False Prompts in SQSK. In this section, we will discuss the situation in MQMK and compare the two scenarios. Compared to SQSK, MQMK has a lower accuracy for the incorrectly selected False Prompt samples. Even when provided with the True Prompt, the performance remains only at 68.34%. This suggests that MQMK has already matched the more recognizable and straightforward samples, and the remaining samples are inherently difficult to recognize. Even with the True Prompt, the performance is limited. The samples selected with the True Prompt in MQMK are easier to match and recognize, achieving an accuracy of 89.01%.
7. Effect of Top-$K$ in Category Aggregation
$K$ $A_{T}$ ($\uparrow$) Matching Rate $(\uparrow$)
@@ 7. Effect of Top-$K$ in Category Aggregation
Table 9. Average accuracy (%) and matching rate (%) under different $K$-category aggregation on 10-task Split Imagenet-R.
In previous experiments, we set $K$ to 1 at all times. Here, we will discuss the performance of aggregating multiple classes and then selecting the prompt. As shown in table 9, when $K$ is 1, both performance and matching rate reach their optimal values. As $K$ increases, both performance and matching rate decline. This suggests that the classes within the same task are not similar in features, and aggregating the matching scores for a group of classes has a negative effect.
8. Dataset Introduction
@@
  • CIFAR-100 is a popular dataset in machine learning, consisting of 100 different classes. It contains 60,000 $\times32$ color images, divided into 50,000 training images and 10,000 test images. Each class contains 600 images, and the categories are diverse, such as animals, vehicles, and household items.
@@
9. Explanation of MQMK-EI and Its Forgetting Behavior
The prompts learned for different tasks are approximately orthogonal. As the model size and the length of the prompts increase, both prompts and features reside in a very high-dimensional space, where random vectors naturally tend to be orthogonal. This approximate orthogonality among different prompts may serve as the foundation for the effectiveness of prompt fusion.
It is a common phenomenon that MQMK-EI exhibits lower forgetting rates but also lower accuracy compared to MQMK, as shown in the Performance Comparison section of the main text. We found that after task $i$ is learned, MQMK achieves higher accuracy than MQMK-EI on the samples from that task. However, as more new tasks are learned, the accuracy of MQMK on task $i$ decreases more significantly, although it still remains higher than that of MQMK-EI, as shown in table 8. This is because the fusion strategy in MQMK-EI introduces interference from previously learned tasks to the current one, leading to performance degradation on the current task. In contrast, MQMK benefits from strong query independence, enabling it to achieve higher performance immediately after training. Nevertheless, as interference accumulates with the learning of more tasks, MQMK eventually experiences performance degradation as well. Based on this analysis, the fewer the number of tasks, the greater the advantage of MQMK.
10. Upper Bound
ImageNet-R CIFAR-100 DomainNet
Perfect Match 84.46 94.91 91.79
Joint Train 79.27 91.79 89.15
Table 10. The comparison between Joint Train and Perfect Match.
The comparison between Joint Train and Perfect Match is shown in table 10. The accuracy of Perfect Match is significantly higher than that of Joint Training. This suggests that the matching approach still has room for improvement and holds considerable promise.